OpenStudy (anonymous):

Find the height of a rectangular prism if the surface area is 148 cm. sq. The width is 4 cm & the length is 5 cm. A)6 B)12 C)7.4 D)14.8

OpenStudy (anonymous):

A) 6.

OpenStudy (anonymous):

@rajathsbhat how did you get that ?

OpenStudy (anonymous):

take the height as x. 

OpenStudy (anonymous):

Area of the bottom+top=20+20=40sq cm

OpenStudy (anonymous):

4x+4x=8x

OpenStudy (anonymous):

5x +5x=10x

OpenStudy (lgbasallote):

\[\large \text{surface area} = lw + lh + wh\] where l = length w = width h = height

OpenStudy (anonymous):

Total area=8x+10x+40=148 18x=108 x=6

OpenStudy (anonymous):

cm

OpenStudy (anonymous):

@rajathsbhat ok got it. thanks !

OpenStudy (anonymous):

yw :)

OpenStudy (lgbasallote):

wait...i think i wrote the formula wrong \[\large \text{surface area} = 2(length \times width) + 2(width \times height) + 2(length \times height)\] sorry for that lol